OpenStudy (anonymous):

is sin x + 1 equivalent to sin2x−sinx−2/sinx−2?

OpenStudy (anonymous):

\[is~\it~\sin^2x~or ~\sin2x\]

OpenStudy (amoodarya):

\[\frac{\sin^2x-sinx-2}{sinx-2}=\frac{(sinx+1)(sinx-2)}{sinx-2}\\sinx-2\neq0\rightarrow \\\frac{(sinx+1)(sinx-2)}{sinx-2}=sinx +1\]